OpenStudy (anonymous):

The length of a rectangle is 1m less than twice its width, and the area of the rectangle is 78m^2. Find the dimensions of the rectangle. Length: ? Width: ?

OpenStudy (anonymous):

length = L width = W L= 2*W-1 L*W= 78 // plug in 2*W-1 for L in the equation (L*W) = 78 solve for W to find the width. post what you get for W.

OpenStudy (anonymous):

I think I got it after that.

OpenStudy (anonymous):

good to hear

OpenStudy (anonymous):

I got -6 and 13/2 but you can't have a negative dimension correct?

OpenStudy (anonymous):

yes, it is impossible to have a negative length. L= 2*W-1 L*W= 78 (2W-1)*(W) = 78 2W^2-W = 78 2W^2-W -78 = 0 quadratic formula. b = -1 a = 2 c = -78 x = [-(-1) +- sqrt((-1)^2-4*(2)(-78))]/(2*2) x = [1 +- sqrt(1+624)]/4 x = [1+- 25]/4 x = (1+25)/4 OR x = (1-25)/4 x = 6.5 OR -6 x = 7.25

OpenStudy (anonymous):

the last x = 7.25 is a typo, it should be 6.5